OpenStudy (cantdoitlikeme):

Final grades are often computed using weighted averages. For example, Fran's teacher uses the following: homework 10%, quizzes 20%, and tests 70%. Fran has an average homework score of 92. Her quiz score average is 68, and her test score average is 81. What final grade can Fran expect? A) 73.5 B) 75.5 C) 78 D) 79.5

OpenStudy (snoopdogg):

homework: 92 * 0.10 = 9.2 quiz: 68 * 0.20 = 13.6 test: 81 * 0.70 = 56.7 9.2 + 13.6 + 56.7 = 79.5 answer: D) 79.5

OpenStudy (snoopdogg):

hope i helped ;)